I am standing in a square garden. the distance between me and the edges of the square garden are 6, 7, 9 and x meters. find x. x is an integer. thanks alot!
Answer by josgarithmetic(39616)

SQUARE garden. This means, the sum of two of those given values or the unknown value is equal to the sum of the other two values or unknown value. You have only very few choices to check.

6+x=7+9?
x could be 10.

6+7=x+9?
x could be 4.

7+x=6+9?
x could be 8.

Only depends on what the size is of the square garden. Edge of 16 meters? 13 meters? 15 meters?